OpenStudy (anonymous):

I have a 8th grade math assignment to do.. Is anyone really good at 8th grade math?

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

no need to fear Mitchal to the recuse!!!!

OpenStudy (anonymous):

im good at 8th grade math, im in 11th grade so yeah XD

OpenStudy (anonymous):

i like that dog

OpenStudy (anonymous):

thanks rottweiler 7 month old

OpenStudy (anonymous):

yea when im older im going to get one

OpenStudy (anonymous):

cool they are awesome dog and very smart too. im training him to be a service dog

OpenStudy (anonymous):

ok sec i will post the questions

OpenStudy (anonymous):

1. What is the value of r in the equation 2(2r + 8) = 3r + 16? (4 points) 0 10 32 2 2. What is the value of b in the equation b + 8 = 13? (4 points) 5 21 8 13 3. Brooke solved the equation below:6r = 726r – 6 = 72 – 6 r = 66Is Brooke's solution correct? If the solution is incorrect, explain why it is incorrect and show the correct steps to solve the equation. (5 points) Font Family Font Size 4. How many solutions can be found for the equation 3x – 2 = 3x + 5? (4 points) Infinitely Many Zero Two One 5. An equation is shown below: 6(2x – 11) + 15 = 3x + 12 Write the steps you will use to solve the equation and explain each step. (5 points) Font Family Font Size 6. The steps below show the incomplete solution to find the value of x for the equation 7x – 3x – 4 = –1 + 21: Step 1: 7x – 3x – 4 = –1 + 21 Step 2: 7x – 3x – 4 = 20 Step 3: 4x – 4 = 20Which of these is most likely the next step? (4 points) 4x = 24 4x = 5 4x = 80 4x = 16 7. What is the value of q in the equation 5q – 10 = 75? (4 points) 65 60 17 85

OpenStudy (anonymous):

@mitchal

OpenStudy (anonymous):

holy crap O_O give me a min XD

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

1. 0 2. 5 3. it is incorrect because there are to many solution. it should always go one solution. 4. one 5. 6(2x – 11) + 15 = 3x + 12 okay you have to distribute the parathesis so 6(2x-11) is 12x - 66 +15= 3x+12 and than you combine like terms the x so you subtract 3 on both side 12x-66+15=3x+12 -3x -3x 9x-66+15=12 +66 +66 9x+15=78 -15 -15 9x=63 /9x /9x x=7 6. 4x=24 7. 17

OpenStudy (anonymous):

Thanks!!!!

OpenStudy (anonymous):

can anyone check that and make sure it is correct?